# A Fisher-Rao gradient flow for entropic mean-field min-max games

### v1

@@ 1.4. Related works
Recent intensive research has been dedicated to examining the convergence of the Wasserstein gradient flow to MNEs within the specific formulation of game (1) in which $F$ is bilinear ($F(\nu,\mu)=\int_{\mathcal{Y}}\int_{\mathcal{X}}f(x,y)\nu(\mathrm{d}x)\mu(\mathrm{d}y)$) and regularized by entropy rather than relative entropy $\operatorname{D_{KL}}.$ The spaces $\mathcal{X}$ and $\mathcal{Y}$ are assumed to be either compact smooth manifolds without boundary, embedded in the Euclidean space, or Euclidean tori, while $f$ exhibits sufficient regularity, typically being at least continuously differentiable with Lipschitz conditions satisfied by $\nabla_{x}f$ and $\nabla_{y}f.$ This line of research is explored in works such as [4, 15, 12, 26].
@@
In this context, [15, 12] investigate the convergence of the Wasserstein gradient flow, proving exponential convergence to the MNE when the players’ flows $(\nu_{t})_{t\geq 0}$ and $(\mu_{t})_{t\geq 0}$ converge at different rates. In [15], the analysis considers the scenario where one player’s flow reaches equilibrium while the other remains governed by the Wasserstein gradient flow equation. Notably, [15, Theorem $5$] shows the convergence (without explicit rate) of the flow $(\nu_{t},\mu_{t})_{t\geq 0}$ to the unique MNE under these separated dynamics.
@@
On the other hand, [12] examines the situation where the players’ flows evolve at varying speeds but with finite timescale separation, meaning that neither player has reached equilibrium. Here, [12, Theorem $2.1$] proves exponential convergence of the finitely timescale-separated Wasserstein gradient flow to the unique MNE, with the convergence rate depending upon regularization and timescale separation parameters. In contrast to [15, 12], we prove that the Fisher-Rao gradient flow (5) converges exponentially fast to the unique MNE while players’ dynamics converge at the same speed.
@@
Other works focused on the convergence of the Wasserstein-Fisher-Rao (WFR) gradient flow, combining both the Wasserstein gradient flow (allowing particles to diffuse in space) and the Fisher-Rao gradient flow (forcing particles to evade low probability regions). Assuming that $F$ is bilinear and $\sigma=0$, [4] investigates the Wasserstein-Fisher-Rao gradient flow’s convergence (without giving explicit rates). For $t_{0}>0$ (dependent on the parameters governing the individual contributions of the Wasserstein and Fisher-Rao components in the WFR gradient flow), and under circumstances where the Fisher-Rao component predominates over the Wasserstein component, [4, Theorem $2$] establishes that the pair $\left(\frac{1}{t_{0}}\int_{0}^{t_{0}}\nu_{s}\mathrm{d}s,\frac{1}{t_{0}}\int_{0}^{t_{0}}\mu_{s}\mathrm{d}s\right)$ is an $\epsilon$-approximate MNE of the game, i.e., $\operatorname{NI}\left(\frac{1}{t_{0}}\int_{0}^{t_{0}}\nu_{s}\mathrm{d}s,\frac{1}{t_{0}}\int_{0}^{t_{0}}\mu_{s}\mathrm{d}s\right)\leq\epsilon,$ for any arbitrarily chosen $\epsilon>0.$
@@
Lastly, [26] introduces a proximal point method that can be viewed as a discrete-time counterpart of the WFR gradient flow. Working within the framework of bilinear $F,$ with $\sigma=0,$ and unique MNE, [26, Theorem $2.2$] establishes the local exponential convergence of the iterates to the unique MNE of the game with respect to the NI error and the WFR distance, provided that the initialization is done in close vicinity to the MNE.
2. Main results
@@
(8) $F(\nu^{\prime},\mu)-F(\nu,\mu)\geq\int_{\mathcal{X}}\frac{\delta F}{\delta\nu}(\nu,\mu,x)(\nu^{\prime}-\nu)(\mathrm{d}x),$
(9) $F(\nu,\mu^{\prime})-F(\nu,\mu)\leq\int_{\mathcal{Y}}\frac{\delta F}{\delta\mu}(\nu,\mu,y)(\mu^{\prime}-\mu)(\mathrm{d}y).$
@@
There exist constants $C_{\nu},C_{\mu}>0$ such that for all $(\nu,\mu)\in\mathcal{P}(\mathcal{X})\times\mathcal{P}(\mathcal{Y})$ and for all $(x,y)\in\mathcal{X}\times\mathcal{Y}$, we have
@@
Suppose $F$ admits second order flat derivatives and that there exist constants $C_{\nu,\nu},C_{\mu,\mu},C_{\nu,\mu},C_{\mu,\nu}>0$ such that for all $(\nu,\mu)\in\mathcal{P}(\mathcal{X})\times\mathcal{P}(\mathcal{Y})$ and for all $(x,y),(x^{\prime},y^{\prime})\in\mathcal{X}\times\mathcal{Y}$, we have
@@
Note that the order of the flat derivatives in $\nu$ and $\mu$ can be interchanged due to [9, Lemma B.2.2.2.2]. Using Assumption 3, it is straightforward to check that there exist constants $C_{\nu}^{\prime},C_{\mu}^{\prime}>0$ such that for all $(\nu,\mu)\in\mathcal{P}_{\text{ac}}(\mathcal{X})\times\mathcal{P}_{\text{ac}}(\mathcal{Y})$, $(\nu^{\prime},\mu^{\prime})\in\mathcal{P}_{\text{ac}}(\mathcal{X})\times\mathcal{P}_{\text{ac}}(\mathcal{Y})$ and all $(x,y)\in\mathcal{X}\times\mathcal{Y}$, we have that
(10) $\left|\frac{\delta F}{\delta\nu}(\nu,\mu,x)-\frac{\delta F}{\delta\nu}(\nu^{\prime},\mu^{\prime},x)\right|\leq C_{\nu}^{\prime}\left(\operatorname{TV}(\nu,\nu^{\prime})+\operatorname{TV}(\mu,\mu^{\prime})\right),$
(11) $\left|\frac{\delta F}{\delta\mu}(\nu,\mu,y)-\frac{\delta F}{\delta\mu}(\nu^{\prime},\mu^{\prime},y)\right|\leq C_{\mu}^{\prime}\left(\operatorname{TV}(\nu,\nu^{\prime})+\operatorname{TV}(\mu,\mu^{\prime})\right).$
@@
An example of a function $F$ which satisfies Assumptions 1, 2, 3 is $F(\nu,\mu)=\int_{\mathcal{Y}}\int_{\mathcal{X}}f(x,y)\nu(\mathrm{d}x)\mu(\mathrm{d}y),$ for a function $f:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$ which is bounded but possibly non-convex-non-concave. Indeed, Assumption 1 is trivially satisfied by such $F,$ while Assumptions 2 and 3 hold due to the boundedness of $f.$ This type of objective function $F$ is prototypical in applications including the training of GANs (see, e.g., [1, 7]) and distributionally robust optimization (see, e.g, [16, 20]).
@@
    (12) $\inf_{x\in\mathcal{X}}\frac{\nu_{0}(x)}{\pi(x)}\geq r_{\nu},\quad\inf_{y\in\mathcal{Y}}\frac{\mu_{0}(y)}{\rho(y)}\geq r_{\mu}.$
  2. (2)
@@
    (13) $\sup_{x\in\mathcal{X}}\frac{\nu_{0}(x)}{\pi(x)}\leq R_{\nu},\quad\sup_{y\in\mathcal{Y}}\frac{\mu_{0}(y)}{\rho(y)}\leq R_{\mu}.$
It can be proved (see, e.g., [9, Proposition A.1.1.1.1]) that the MNE $(\nu_{\sigma}^{*},\mu_{\sigma}^{*})$ of 1 is given implicitly by the equations
(14) $\nu_{\sigma}^{*}(x)=\frac{1}{Z(\nu_{\sigma}^{*},\mu_{\sigma}^{*})}\exp{\left(-\frac{2}{\sigma^{2}}\frac{\delta F}{\delta\nu}(\nu_{\sigma}^{*},\mu_{\sigma}^{*},x)-U^{\pi}(x)\right)},$
(15) $\mu_{\sigma}^{*}(y)=\frac{1}{Z^{\prime}(\nu_{\sigma}^{*},\mu_{\sigma}^{*})}\exp{\left(\frac{2}{\sigma^{2}}\frac{\delta F}{\delta\mu}(\nu_{\sigma}^{*},\mu_{\sigma}^{*},y)-U^{\rho}(y)\right)},$
@@
(16) $\bar{r}_{\nu}\leq\frac{\nu_{0}(x)}{\nu_{\sigma}^{*}(x)}\leq\bar{R}_{\nu},\quad\bar{r}_{\mu}\leq\frac{\mu_{0}(y)}{\mu_{\sigma}^{*}(y)}\leq\bar{R}_{\mu}.$
We emphasize that Assumption 4 is natural in the context of Fisher-Rao flows. From (5), we observe that the support of the measures along the gradient flow does not increase. Thus, it is essential that the initialization is comparable with the MNE (cf. (16)) as reflected in Assumption 4. It is observed in [11], that Assumption 4 can be understood as a “warm start” type of condition.
@@
(17) $\begin{cases}\partial_{t}\nu_{t}(x)=-a(\nu_{t},\mu_{t},x)\nu_{t}(x),\\ \partial_{t}\mu_{t}(y)=b(\nu_{t},\mu_{t},y)\mu_{t}(y),\end{cases}$
with initial condition $(\nu_{0},\mu_{0})\in\mathcal{P}_{\text{ac}}(\mathcal{X})\times\mathcal{P}_{\text{ac}}(\mathcal{Y})$.
The following result extends [11, Theorem $2.1$] to the case of two-player zero-sum games by showing that the Fisher-Rao gradient flow (17) admits a unique solution $(\nu_{t},\mu_{t})_{t\geq 0}.$
@@
Suppose that Assumption 2, 3 and condition (13) from Assumption 4 hold. Then the system of equations (17) has a unique solution $(\nu_{t},\mu_{t})_{t\geq 0}$. Moreover, for $t\geq 0$,
(18) $\operatorname{D_{KL}}(\nu_{t}|\pi)\leq 2\log R_{\nu}+\frac{4}{\sigma^{2}}C_{\nu},\quad\operatorname{D_{KL}}(\mu_{t}|\rho)\leq 2\log R_{\mu}+\frac{4}{\sigma^{2}}C_{\mu}$
and there exist constants $R_{1,\nu},R_{1,\mu}>1$ such that for all $t\geq 0$,
(19) $\sup_{x\in\mathcal{X}}\frac{\nu_{t}(x)}{\pi(x)}\leq R_{1,\nu},\quad\sup_{y\in\mathcal{Y}}\frac{\mu_{t}(y)}{\rho(y)}\leq R_{1,\mu}.$
Additionally, if condition (12) from Assumption 4 holds, then there exist constants $r_{1,\nu}$, $r_{1,\mu}>0$ such that for all $t\geq 0$,
(20) $\inf_{x\in\mathcal{X}}\frac{\nu_{t}(x)}{\pi(x)}\geq r_{1,\nu},\quad\inf_{y\in\mathcal{Y}}\frac{\mu_{t}(y)}{\rho(y)}\geq r_{1,\mu}.$
@@
Next, we state the other main result of this paper. We prove two different types of convergence results for the FR dynamics (17) in min-max games given by (1), one in terms of the players’ strategies and the other in terms of the payoff function. Whereas the proof of the existence of the flow (Theorem 2.2) follows a route similar to [11], the proof of convergence in Theorem 2.3 is significantly different from [11] since, as indicated in the introduction, in the present paper convergence has to be studied with respect to appropriately chosen Lyapunov functions and, moreover, we do not rely on the Polyak-Łojasiewicz inequality.
@@
Suppose that Assumption 2, 3 and 4 hold and let $(\nu,\mu)\in\mathcal{P}_{\text{ac}}(\mathcal{X})\times\mathcal{P}_{\text{ac}}(\mathcal{Y}).$ Then the map $t\mapsto\operatorname{D_{KL}}(\nu|\nu_{t})+\operatorname{D_{KL}}(\mu|\mu_{t})$ is differentiable along the FR dynamics (17). Suppose furthermore that Assumption 1 holds and that $(\nu_{0},\mu_{0})$ are chosen such that $\max_{\nu\in\mathcal{P}_{\text{ac}}(\mathcal{X})}\operatorname{D_{KL}}(\nu|\nu_{0})+\max_{\mu\in\mathcal{P}_{\text{ac}}(\mathcal{Y})}\operatorname{D_{KL}}(\mu|\mu_{0})<\infty.$ Then, for all $t\geq 0$, we have
@@
(21) $\operatorname{NI}\left(\frac{1}{t}\int_{0}^{t}\nu_{s}\mathrm{d}s,\frac{1}{t}\int_{0}^{t}\mu_{s}\mathrm{d}s\right)\leq\frac{1}{t}\left(\max_{\nu\in\mathcal{P}_{\text{ac}}(\mathcal{X})}\operatorname{D_{KL}}(\nu|\nu_{0})+\max_{\mu\in\mathcal{P}_{\text{ac}}(\mathcal{Y})}\operatorname{D_{KL}}(\mu|\mu_{0})\right).$
In game theoretic language, the first result of Theorem 2.3 says that convergence of the FR dynamics (17) to the unique MNE $(\nu_{\sigma}^{*},\mu_{\sigma}^{*})$ in terms of the strategies $\nu_{t}$ and $\mu_{t}$ is achieved with exponential rate depending on $\sigma.$ On the other hand, the second result gives a convergence estimate in terms of the NI error, which in turn is expressed as a function of the payoff $V^{\sigma}.$ More precisely, the NI error of the time-averaged flow converges to 00 with rate $\mathcal{O}\left(1/t\right)$ regardless of the value of $\sigma.$

### v2

@@ 1.4 Related works
Recent intensive research has been dedicated to examining the convergence of the Wasserstein gradient flow to MNEs within the specific formulation of game (1) in which $F$ is bilinear ($F(\nu,\mu)=\int_{\mathcal{Y}}\int_{\mathcal{X}}f(x,y)\nu(\mathrm{d}x)\mu(\mathrm{d}y)$) and regularized by entropy rather than relative entropy $\operatorname{D_{KL}}.$ The spaces $\mathcal{X}$ and $\mathcal{Y}$ are assumed to be either compact smooth manifolds without boundary, embedded in the Euclidean space, or Euclidean tori, while $f$ exhibits sufficient regularity, typically being at least continuously differentiable with Lipschitz conditions satisfied by $\nabla_{x}f$ and $\nabla_{y}f.$ This line of research is explored in works such as Domingo-Enrich et al. (2020); Ma & Ying (2022); Lu (2023); Wang & Chizat (2023).
@@
In this context, Ma & Ying (2022); Lu (2023) investigate the convergence of the Wasserstein gradient flow, proving exponential convergence to the MNE when the players’ flows $(\nu_{t})_{t\geq 0}$ and $(\mu_{t})_{t\geq 0}$ converge at different rates. In Ma & Ying (2022), the analysis considers the scenario where one player’s flow reaches equilibrium while the other remains governed by the Wasserstein gradient flow equation. Notably, Ma & Ying (2022, Theorem $5$) shows the convergence (without explicit rate) of the flow $(\nu_{t},\mu_{t})_{t\geq 0}$ to the unique MNE under these separated dynamics.
@@
On the other hand, Lu (2023) examines the situation where the players’ flows evolve at varying speeds but with finite timescale separation, meaning that neither player has reached equilibrium. Here, Lu (2023, Theorem $2.1$) proves exponential convergence of the finitely timescale-separated Wasserstein gradient flow to the unique MNE, with the convergence rate depending upon regularization and timescale separation parameters. In contrast to Ma & Ying (2022); Lu (2023), we prove that the Fisher-Rao gradient flow (5) converges exponentially fast to the unique MNE while players’ dynamics converge at the same speed.
@@
Other works such as (Domingo-Enrich et al., 2020; Wang & Chizat, 2023) focused on the convergence of the Wasserstein-Fisher-Rao (WFR) gradient flow, combining both the Wasserstein gradient flow (allowing particles to diffuse in space) and the Fisher-Rao gradient flow (forcing particles to evade low probability regions). Assuming that $F$ is bilinear and $\sigma=0$, Domingo-Enrich et al. (2020) investigates the Wasserstein-Fisher-Rao gradient flow’s convergence (without giving explicit rates). For $t_{0}>0$ (dependent on the parameters governing the individual contributions of the Wasserstein and Fisher-Rao components in the WFR gradient flow), and under circumstances where the Fisher-Rao component predominates over the Wasserstein component, Domingo-Enrich et al. (2020) prove a characterization result of $\epsilon$-approximate MNE for the WFR gradient flow. More precisely, Domingo-Enrich et al. (2020, Theorem $2$) establishes that the pair $\left(\frac{1}{t_{0}}\int_{0}^{t_{0}}\nu_{s}\mathrm{d}s,\frac{1}{t_{0}}\int_{0}^{t_{0}}\mu_{s}\mathrm{d}s\right)$ is an $\epsilon$-approximate MNE of the game, i.e., $\operatorname{NI}\left(\frac{1}{t_{0}}\int_{0}^{t_{0}}\nu_{s}\mathrm{d}s,\frac{1}{t_{0}}\int_{0}^{t_{0}}\mu_{s}\mathrm{d}s\right)\leq\epsilon,$ for any arbitrarily chosen $\epsilon>0.$
@@
Lastly, Wang & Chizat (2023) introduces a proximal point method that can be viewed as a discrete-time counterpart of the WFR gradient flow. Working within the framework of bilinear $F,$ with $\sigma=0,$ and unique MNE, Wang & Chizat (2023, Theorem $2.2$) establishes the local exponential convergence of the iterates to the unique MNE of the game with respect to the NI error and the WFR distance, provided that the initialization is done in close vicinity to the MNE. The algorithm proposed by Wang & Chizat (2023) assumes that both players update the positions and weights of the particles simultaneously. However, this is not the only possible discretization for the WFR gradient flow. As highlighted in Lascu et al. (2024) for the discrete-time Fisher-Rao gradient flow, the convergence rates of the dynamics in a game differ depending on the order the players move: either simultaneously (players move at the same time) or sequentially (each player moves upon observing the opponents’ moves). Although the continuous-time analysis of the gradient flow does not capture these two situations, it can serve as a guide for designing discrete-time approximations of implementable algorithms.
Moreover, it is argued in Wang & Chizat (2023) that the results of Theorem $2.2$ hold only for discrete-time algorithms, and do not imply the convergence of the continuous-time WFR gradient flow because sending the step-sizes of the time discretization to zero will force the initialization $(\nu_{0},\mu_{0})$ to be already an MNE (see the discussion after Theorem $2.2$). Unlike the assumption in Wang & Chizat (2023) that guarantees convergence of the discrete-time algorithm as long as the initialization is close to the MNE, i.e., $\operatorname{NI}(\nu_{0},\mu_{0})\leq r_{0},$ for some $r_{0}>0,$ depending on the step-sizes, our warm start condition (Assumption 4) for the Fisher-Rao gradient flow only imposes absolute continuity and comparability with a priori known reference measures $\pi$ and $\rho.$
Furthermore, Wang & Chizat (2023) discuss that if $(\nu_{0},\mu_{0})$ is supported on the entire space $\mathcal{X}$ and $\mathcal{Y},$ respectively, then it is expected that the convergence of the discrete-time Fisher-Rao dynamics to an MNE occurs at sub-linear rate in the worst case, and not an exponential rate (see Appendix D). Also, if the strategy spaces $\mathcal{X}$ and $\mathcal{Y}$ are finite and the initialization $(\nu_{0},\mu_{0})$ is supported on a large number of points uniformly covering $\mathcal{X}$ and $\mathcal{Y},$ respectively, there is no guarantee for last-iterate convergence of discrete-time Fisher-Rao dynamics to an MNE of $V^{0}.$ As showed by Wei et al. (2021); Daskalakis & Panageas (2019), last-iterate convergence guarantees for games on finite strategy spaces $\mathcal{X}$ and $\mathcal{Y}$ hold under the assumption that the MNE is unique, which may not hold for $V^{0}.$
While it seems natural to combine the Wasserstein and the Fisher-Rao gradient flows, rigorously proving explicit convergence rates for the continuous-time WFR gradient flow for games such as (1) is a challenging and, to our knowledge, still open problem. In this work, we provide an additional step by establishing last-iterate exponential convergence of the Fisher-Rao gradient flow to the unique MNE measured in both KL divergence and NI error (see Theorem 2.3).
2 Main results
@@
$F(\nu^{\prime},\mu)-F(\nu,\mu)\geq\int_{\mathcal{X}}\frac{\delta F}{\delta\nu}(\nu,\mu,x)(\nu^{\prime}-\nu)(\mathrm{d}x),$ (10)
$F(\nu,\mu^{\prime})-F(\nu,\mu)\leq\int_{\mathcal{Y}}\frac{\delta F}{\delta\mu}(\nu,\mu,y)(\mu^{\prime}-\mu)(\mathrm{d}y).$ (11)
@@
Suppose $F$ admits first order flat derivatives with respect to both $\nu$ and $\mu$ as stated in Definition B.1 and there exist constants $C_{\nu},C_{\mu}>0$ such that for all $(\nu,\mu)\in\mathcal{P}(\mathcal{X})\times\mathcal{P}(\mathcal{Y})$ and for all $(x,y)\in\mathcal{X}\times\mathcal{Y}$, we have
@@
Suppose $F$ admits second order flat derivatives as stated in Definition B.1 and there exist constants $C_{\nu,\nu},C_{\mu,\mu},C_{\nu,\mu},C_{\mu,\nu}>0$ such that for all $(\nu,\mu)\in\mathcal{P}(\mathcal{X})\times\mathcal{P}(\mathcal{Y})$ and for all $(x,y),(x^{\prime},y^{\prime})\in\mathcal{X}\times\mathcal{Y}$, we have
@@
Note that the order of the flat derivatives in $\nu$ and $\mu$ can be interchanged due to Lascu et al. (2023, Lemma B.2.2.2.2). Using Assumption 3, it is straightforward to check that there exist constants $C_{\nu}^{\prime},C_{\mu}^{\prime}>0$ such that for all $(\nu,\mu)\in\mathcal{P}_{\text{ac}}(\mathcal{X})\times\mathcal{P}_{\text{ac}}(\mathcal{Y})$, $(\nu^{\prime},\mu^{\prime})\in\mathcal{P}_{\text{ac}}(\mathcal{X})\times\mathcal{P}_{\text{ac}}(\mathcal{Y})$ and all $(x,y)\in\mathcal{X}\times\mathcal{Y}$, we have that
$\left|\frac{\delta F}{\delta\nu}(\nu,\mu,x)-\frac{\delta F}{\delta\nu}(\nu^{\prime},\mu^{\prime},x)\right|\leq C_{\nu}^{\prime}\left(\operatorname{TV}(\nu,\nu^{\prime})+\operatorname{TV}(\mu,\mu^{\prime})\right),$ (12)
$\left|\frac{\delta F}{\delta\mu}(\nu,\mu,y)-\frac{\delta F}{\delta\mu}(\nu^{\prime},\mu^{\prime},y)\right|\leq C_{\mu}^{\prime}\left(\operatorname{TV}(\nu,\nu^{\prime})+\operatorname{TV}(\mu,\mu^{\prime})\right).$ (13)
@@
An example of a function $F$ which satisfies Assumptions 1, 2, 3 is $F(\nu,\mu)=\int_{\mathcal{Y}}\int_{\mathcal{X}}f(x,y)\nu(\mathrm{d}x)\mu(\mathrm{d}y),$ for a function $f:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$ which is bounded but possibly non-convex-non-concave. Indeed, Assumption 1 is trivially satisfied by such $F,$ while Assumptions 2 and 3 hold due to the boundedness of $f.$ This type of objective function $F$ is prototypical in applications including the training of GANs (see, e.g., Arjovsky et al. (2017); Hsieh et al. (2019)) and distributionally robust optimization (see, e.g, Madry et al. (2018); Sinha et al. (2018)).
Another example that can be viewed as a particular case of our general framework is the objective function of Wasserstein-GANs (WGANs) with gradient penalty (Gulrajani et al., 2017; Petzka et al., 2018). Following Conforti et al. (2023); Kazeykina et al. (2022), GANs can be framed as a min-max problem over the space of probability measures. Consider the following class of parameterized functions as the class of choices for the discriminator:
$\{x\mapsto\mathbb{E}[f(Y,x)]:\operatorname{Law}(Y)=\mu\in\mathcal{P}(\mathcal{Y})\}.$
The map $x\mapsto\mathbb{E}_{Y\sim\mu}[f(Y,x)]$ is the output of a two-layer neural network with parameters $Y$ and bounded, continuous and non-constant activation function $f.$ The objective function of the GAN reads
$\min_{\nu\in\mathcal{P}(\mathcal{X})}\max_{\mu\in\mathcal{P}(\mathcal{Y})}\int_{\mathcal{X}}\mathbb{E}_{Y\sim\mu}[f(Y,x)](\nu-\bar{\nu})(\mathrm{d}x)=\min_{\nu\in\mathcal{P}(\mathcal{X})}\max_{\mu\in\mathcal{P}(\mathcal{Y})}\int_{\mathcal{Y}}\int_{\mathcal{X}}f(x,y)(\nu-\bar{\nu})(\mathrm{d}x)\mu(\mathrm{d}y),$
where $\bar{\nu}$ is the distribution of the true data. Assuming that $\mathcal{X}\ni x\mapsto f(x,y)\in\mathbb{R}$ is differentiable and $1$-Lipschitz, we retrieve the WGAN. The gradient penalty regularization terms proposed by Gulrajani et al. (2017); Petzka et al. (2018) are added in order to ensure that the discriminator remains $1$-Lipschitz along the interpolation between the generated and true data. Thus, the objective function can be expressed as
$\min_{\nu\in\mathcal{P}(\mathcal{X})}\max_{\mu\in\mathcal{P}(\mathcal{Y})}\Bigg{\{}\int_{\mathcal{Y}}\int_{\mathcal{X}}f(x,y)(\nu-\bar{\nu})(\mathrm{d}x)\mu(\mathrm{d}y)+\lambda t\int_{\mathcal{Y}}\int_{\mathcal{X}}\left(\|\nabla_{x}f(x,y)\|-1\right)^{2}(\nu-\bar{\nu})(\mathrm{d}x)\mu(\mathrm{d}y)$
$+\lambda(1-t)g\left(\int_{\mathcal{Y}}\int_{\mathcal{X}}\left(\|\nabla_{x}f(x,y)\|-1\right)^{2}\bar{\nu}(\mathrm{d}x)\mu(\mathrm{d}y)\right)\Bigg{\}},$
where $g:\mathbb{R}\to\mathbb{R}$ is a twice differentiable concave function with bounded derivatives, $\lambda>0$ is a regularization parameter and $t\in[0,1].$ Since $g$ is concave, the term inside $g$ is linear in $\mu$ and independent of $\nu,$ it follows that the third term in the optimization problem is concave in $\mu.$ This fact together with the linearity in $\nu$ and $\mu$ of the first two terms shows that the objective function satisfies Assumption 1. Since $f$ is bounded, differentiable and $1$-Lipschitz, and $g$ is twice differentiable with bounded derivatives, Assumptions 2 and 3 are also satisfied.
@@
    $\inf_{x\in\mathcal{X}}\frac{\nu_{0}(x)}{\pi(x)}\geq r_{\nu},\quad\inf_{y\in\mathcal{Y}}\frac{\mu_{0}(y)}{\rho(y)}\geq r_{\mu}.$ (14)
  2. 2.
@@
    $\sup_{x\in\mathcal{X}}\frac{\nu_{0}(x)}{\pi(x)}\leq R_{\nu},\quad\sup_{y\in\mathcal{Y}}\frac{\mu_{0}(y)}{\rho(y)}\leq R_{\mu}.$ (15)
It can be proved (see, e.g., Lascu et al. (2023, Proposition A.1.1.1.1)) that the MNE $(\nu_{\sigma}^{*},\mu_{\sigma}^{*})$ of (1) satisfies the fixed point equations
$\nu_{\sigma}^{*}(x)=\frac{1}{Z(\nu_{\sigma}^{*},\mu_{\sigma}^{*})}\exp{\left(-\frac{2}{\sigma^{2}}\frac{\delta F}{\delta\nu}(\nu_{\sigma}^{*},\mu_{\sigma}^{*},x)-U^{\pi}(x)\right)},$ (16)
$\mu_{\sigma}^{*}(y)=\frac{1}{Z^{\prime}(\nu_{\sigma}^{*},\mu_{\sigma}^{*})}\exp{\left(\frac{2}{\sigma^{2}}\frac{\delta F}{\delta\mu}(\nu_{\sigma}^{*},\mu_{\sigma}^{*},y)-U^{\rho}(y)\right)},$ (17)
@@
$\bar{r}_{\nu}\leq\frac{\nu_{0}(x)}{\nu_{\sigma}^{*}(x)}\leq\bar{R}_{\nu},\quad\bar{r}_{\mu}\leq\frac{\mu_{0}(y)}{\mu_{\sigma}^{*}(y)}\leq\bar{R}_{\mu}.$ (18)
We emphasize that Assumption 4 is natural in the context of Fisher-Rao flows. From (5), we observe that the support of the measures along the gradient flow does not increase. Thus, it is essential that the initialization is comparable with the MNE (cf. (18)) as reflected in Assumption 4. It is observed in Liu et al. (2023), that Assumption 4 can be understood as a “warm start” type of condition.
@@
$\begin{cases}\partial_{t}\nu_{t}(x)=-a(\nu_{t},\mu_{t},x)\nu_{t}(x),\\ \partial_{t}\mu_{t}(y)=b(\nu_{t},\mu_{t},y)\mu_{t}(y),\end{cases}$ (19)
with initial condition $(\nu_{0},\mu_{0})\in\mathcal{P}_{\text{ac}}(\mathcal{X})\times\mathcal{P}_{\text{ac}}(\mathcal{Y})$. We will establish the existence of a solution to (19) in Theorem 2.2. As we will demonstrate, the flow of densities $(\nu_{t},\mu_{t})_{t\geq 0}$ is differentiable in time (cf. equation (35)), and hence the solution to (19) can be interpreted just as a classical solution to an ordinary differential equation.
The following result extends Liu et al. (2023, Theorem $2.1$) to the case of two-player zero-sum games by showing that the Fisher-Rao gradient flow (19) admits a unique solution $(\nu_{t},\mu_{t})_{t\geq 0}.$
@@
Suppose that Assumption 2, 3 and condition (15) from Assumption 4 hold. Then for each $(\nu_{0},\mu_{0})\in\mathcal{P}_{\text{ac}}(\mathcal{X})\times\mathcal{P}_{\text{ac}}(\mathcal{Y}),$ there exists a unique pair of continuous and differentiable in time flows $(\nu_{t},\mu_{t})_{t\in[0,\infty)}\in\mathcal{P}_{\text{ac}}(\mathcal{X})\times\mathcal{P}_{\text{ac}}(\mathcal{Y})$ satisfying the system of equations (19). Moreover, for $t>0$,
$\operatorname{D_{KL}}(\nu_{t}|\pi)\leq 2\log R_{\nu}+\frac{4}{\sigma^{2}}C_{\nu},\quad\operatorname{D_{KL}}(\mu_{t}|\rho)\leq 2\log R_{\mu}+\frac{4}{\sigma^{2}}C_{\mu}$ (20)
and there exist constants $R_{1,\nu},R_{1,\mu}>1$ such that for all $t>0$,
$\sup_{x\in\mathcal{X}}\frac{\nu_{t}(x)}{\pi(x)}\leq R_{1,\nu},\quad\sup_{y\in\mathcal{Y}}\frac{\mu_{t}(y)}{\rho(y)}\leq R_{1,\mu}.$ (21)
Additionally, if condition (14) from Assumption 4 holds, then there exist constants $r_{1,\nu}$, $r_{1,\mu}>0$ such that for all $t>0$,
$\inf_{x\in\mathcal{X}}\frac{\nu_{t}(x)}{\pi(x)}\geq r_{1,\nu},\quad\inf_{y\in\mathcal{Y}}\frac{\mu_{t}(y)}{\rho(y)}\geq r_{1,\mu}.$ (22)
@@
Next, we state the other main result of this paper. We prove two different types of convergence results for the FR dynamics (19) in min-max games given by (1), one in terms of the players’ strategies and the other in terms of the payoff function. Whereas the proof of the existence of the flow (Theorem 2.2) follows a route similar to Liu et al. (2023), the proof of convergence in Theorem 2.3 is significantly different from Liu et al. (2023) since, as indicated in the introduction, in the present paper, convergence has to be studied with respect to appropriately chosen Lyapunov functions and, moreover, we do not rely on the Polyak-Łojasiewicz inequality.
@@
Suppose that Assumption 2, 3 and 4 hold and let $(\nu,\mu)\in\mathcal{P}_{\text{ac}}(\mathcal{X})\times\mathcal{P}_{\text{ac}}(\mathcal{Y}).$ Then the map $t\mapsto\operatorname{D_{KL}}(\nu|\nu_{t})+\operatorname{D_{KL}}(\mu|\mu_{t})$ is differentiable along the FR dynamics (19). Suppose furthermore that Assumption 1 holds. Then, for all $t>0$, we have
@@
$\operatorname{NI}(\nu_{t},\mu_{t})\leq 2C_{\sigma}e^{-\frac{\sigma^{2}}{4}t}\sqrt{\operatorname{D_{KL}}(\nu_{\sigma}^{*}|\nu_{0})+\operatorname{D_{KL}}(\mu_{\sigma}^{*}|\mu_{0})}.$ (23)
In game theoretic language, the first result of Theorem 2.3 says that convergence of the FR dynamics (19) to the unique MNE $(\nu_{\sigma}^{*},\mu_{\sigma}^{*})$ in terms of the strategies $\nu_{t}$ and $\mu_{t}$ is achieved with exponential rate depending on $\sigma,$ while the second result shows exponential convergence in terms of the payoff function $V^{\sigma}.$